# Software Performance Engineering for Foundation Model-Powered Software (FMware)

### v1

@@ V Conclusion
This paper has explored the emerging SPE challenges for FMware, highlighting numerous opportunities for innovation to enhance current practices. We explored four major challenges spanning the software lifecycle for FMware and discussed our attempt to addressing SLA-aware multi-tenant serving. Our vision of FMware Runtime marks only the first step towards tackling these challenges, stemming from our experience working with FMware developers to resolve performance issues, discussions with world-renowned scholars, and comprehensive surveys. The insights presented here aim to help developers address performance concerns more effectively than traditional SPE methodologies. We encourage both researchers and practitioners in the SPE community to advancing FMware performance engineering. The unique challenges outlined in this paper represent critical areas for future work, as the SPE field continues to evolve alongside next-generation AI-powered software.
VI Disclaimer
@@
Any opinions, findings, conclusions, or recommendations expressed in this material are those of the author(s) and do not reflect the views of Huawei. Also, ChatGPT-4.0 was used for copy-editing. All experiments, analysis, writing, and results were performed by the authors, who also thoroughly reviewed the final content. This complies with IEEE and ACM policies on AI use in publications.

### v2

@@ 5. Conclusion
This paper has explored the emerging SPE challenges for FMware, highlighting numerous opportunities for innovation to enhance current practices. We explored four major challenges spanning the software lifecycle for FMware and discussed our attempt to addressing SLA-aware multi-tenant serving. Our vision of FMware Runtime marks only the first step towards tackling these challenges, stemming from our experience working with FMware developers to resolve performance issues, discussions with world-renowned scholars, and comprehensive surveys. The insights presented here aim to help developers address performance concerns more effectively than traditional SPE methodologies. Future work should focus on establishing benchmarks and tooling to enable researchers and practitioners to systematically observe real-world FMware performance issues, empirically validate these four SPE challenges, and optimize performance across a broader set of Promptware and Agentware projects. We encourage both researchers and practitioners in the SPE community to advancing FMware performance engineering. The unique challenges outlined in this paper represent critical areas for future work, as the SPE field continues to evolve alongside next-generation AI-powered software.
6. Disclaimer
@@
Any opinions, findings, conclusions, or recommendations expressed in this material are those of the author(s) and do not reflect the views of Huawei. Also, ChatGPT was used for copy-editing. All experiments, analysis, writing, and results were performed by the authors, who also thoroughly reviewed the final content. This complies with IEEE and ACM policies on AI use in publications.